-8x-8y=-16 -- [find two points on this line]
Answer by TimothyLamb(4379)

Get equation into slope-intercept form: y = mx + b
-8x -8y = -16
-8x + 16 = 8y
y = -x + 2
---
p1:
try x = 0
y = -x + 2
y = 2
p1 = (0, 2)
---
p2:
try x = 1
y = -x + 2
y = -1 + 2
y = 1
p2 = (1, 1)
---
Answer:
Draw a line between the points p1 and p2 given above.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php